Anita can rent a car for a day from Smooth Rent A Car for $60 plus $0.75/km or from Zippy Car Rental for $ 1.25/km.
A) write a system linear equations representing the cost to rent the car from each company for a day
B) calculate the number of km for which the rental charges from either company will be the same. Show all of your work.

Answer by addingup(3677)

60+0.75x = Smooth
1.25x = Zippy
--------------------------
60+0.75x = 1.25x
-0.50x = -60
x = -60/-0.50 use your calculator, divide. Remember -/- = + so you can ignore the -sign, your answer will be a positive number anyway.